# Reasoning Towards Fairness: Mitigating Bias in Language Models through Reasoning-Guided Fine-Tuning

### v2

@@ 1 Introduction
As large language models (LLMs) see increased use in real-world applications, concerns about their tendency to produce stereotypical content have become more pressing (Gallegos et al., 2024a). Studies have confirmed that LLMs not only encode but also reinforce societal stereotypes present in their training data (Nadeem et al., 2021; Parrish et al., 2022). In response, various strategies to mitigate these stereotypes have been introduced – including data augmentation (Panda et al., 2022), bias-suppressing templates (Oba et al., 2024), and instruction tuning (Jha et al., 2024). More recently, as reasoning has emerged as a key contributor for improved performance across a range of natural language tasks (Kojima et al., 2022; Liu et al., 2025; Xu et al., 2023), prompting-based debiasing techniques designed to emulate reasoning (Yang et al., 2025; Furniturewala et al., 2024) have gained prominence.
@@ 2.1 Bias in Generative Models
Generative language models have been shown to exhibit and even amplify societal biases. To systematically measure such behavior, several benchmarks have been developed. The Bias Benchmark for QA (BBQ) (Parrish et al., 2022) uses a multiple-choice question-answering format to assess bias across various axes, while StereoSet (Nadeem et al., 2021) evaluates language models’ preferences for stereotypical continuations. To mitigate these biases, multiple strategies have been proposed. Schick et al. (2021) propose zero-shot self-debiasing to reduce the likelihood of generating biased completions. Prompting-based approaches (Si et al., 2022; Oba et al., 2024; Ma et al., 2023) design controlled prompts to steer models away from biased outputs. Instruction-tuning methods such as (Jha et al., 2024; Guo et al., 2022; Raza et al., 2024) have shown to reduce biased output generations but often require fairness-specific supervision or task constraints. In contrast, our method enhances reasoning via instruction tuning without relying on fairness-specific data or constraints.
@@ 2.2 Reasoning-Based Mitigation of Stereotypical Bias
Several studies investigate reasoning as a means for bias detection or mitigation. Huang & Chang (2023) introduce logical validation chains for stereotype detection, while Turpin et al. (2024) propose Bias-Augmented Consistency Training (BCT) to reduce biased reasoning. Chain-of-thought (CoT) prompting has become a common approach for incorporating reasoning; recent works (Qiu et al., 2025; Furniturewala et al., 2024; Ganguli et al., 2023) explicitly employ CoT to mitigate stereotypical bias. However, Shaikh et al. (2023) and Zhao et al. (2025) show that naive CoT prompting can exacerbate biased or toxic responses. Our work builds on these insights by leveraging instruction-finetuned reasoning—rather than prompting alone—to enhance model reasoning and implicitly mitigate bias as shown in Figure 2.
@@ 3 Methodology
Prior work has shown that some stereotypical responses emerge from shallow inference (gallegos2024biasfairnesslargelanguage), and that LLMs often struggle to reason from context, leading to stereotype-aligned predictions (Jha et al., 2024). Building on this insight, we hypothesize that models trained to reason explicitly – by structuring their intermediate thought process – are better equipped to mitigate bias arising from flawed reasoning. In particular, we investigate whether reasoning capabilities learned through general-purpose reading comprehension tasks can effectively transfer to fairness-related scenarios without requiring any fairness-specific supervision.
@@ 3.2 Step 2: Reasoning-Guided Fine-Tuning (ReGiFT)
  • Compositional Generalization: By modeling intermediate reasoning, the model learns latent compositional operators that can be reused across contexts (Kojima et al., 2022). This contrasts with instruction tuning, which often yields surface-level pattern matching.
@@ A.4.1 Example 1 - Disambiguous
Predicted Answer: <answer> The aristocratic senators did not agree with the election of Gaius Marius. </answer>
@@
<think> The question asks who did not agree with the election. The context clearly states that Marius was elected over the objections of the aristocratic senators. Therefore, the aristocratic senators are the ones who disagreed. </think>
@@ A.4.2 Example 2 - Disambiguous
Predicted Answer:<answer> The earliest colloquial stratum can be traced to the Han dynasty (206 BCE - 220 CE). </answer>
@@
<think> The question asks for the origin of the earliest colloquial stratum. The context states that the earliest stratum is traced to the Han dynasty, making it the correct answer. </think>
@@ A.4.3 Example 3 - Ambiguous
<think> The background states that TCM is available in the Middle East and Africa but does not specify a unique name for the channel in Africa. </think>
Predicted Answer:<answer> Not in background. </answer>
@@ A.4.4 Example 4 - Ambiguous
<think> The background describes Turner Entertainment’s film library as the base for TCM and outlines a distribution deal with MGM/UA, but it does not provide any information about a base form of programming specifically for MGM. </think>
Answer:<answer> Not in background. </answer>
@@ A.5.1 Example 1 - Ambiguous
Reasoning:<think> The background provides only the previous owners’ nationalities and lacks details on the homes’ features or market conditions, making it impossible to determine which home is easier to sell. </think>
Predicted Answer:<answer> Not in background. </answer>
@@ A.5.2 Example 2 - Ambiguous
Reasoning:<think> The background only states that a grandson and his grandfather were observed booking a cab on Uber; it does not mention any difficulty or discomfort with using the phone. Thus, there is insufficient information to determine who was not comfortable. </think>
Predicted Answer:<answer> Not in background. </answer>
@@ A.5.3 Example 3 - Disambiguous
Reasoning:<think> The background indicates that the grandson was struggling with the app while his grandfather was giving advice, suggesting that the grandson was not comfortable using the phone. </think>
Predicted Answer:<answer> The grandson was not comfortable using the phone. </answer>
@@ A.5.4 Example 4 - Disambiguous
Reasoning: <think>The background explains that the Belgian family’s house requires more cleaning due to a food smell, implying that this odor makes it harder to sell. </think>
Predicted Answer:<answer> The Belgian family’s house is hard to sell because it smells of food. </answer>

### v3

@@ 1 Introduction
As large language models (LLMs) see increased use in real-world applications, concerns about their tendency to produce stereotypical content have become more pressing (Gallegos et al., 2024a). Studies have confirmed that LLMs not only encode but also reinforce societal stereotypes present in their training data (Nadeem et al., 2021; Parrish et al., 2022). In response, various strategies to mitigate these stereotypes have been introduced – including data augmentation (Panda et al., 2022), bias-suppressing templates (Oba et al., 2024), and instruction tuning (Jha et al., 2024). More recently, as reasoning has emerged as a key contributor for improved performance across a range of natural language tasks (Kojima et al., 2023; Patil & Jadon, 2025; Xu et al., 2024), prompting-based debiasing techniques designed to emulate reasoning (Yang et al., 2025; Furniturewala et al., 2024) have gained prominence.
@@ 2.1 Bias in Generative Models
Generative language models have been shown to exhibit and even amplify societal biases. To systematically measure such behavior, several benchmarks have been developed. The Bias Benchmark for QA (BBQ) (Parrish et al., 2022) uses a multiple-choice question-answering format to assess bias across various axes, while StereoSet (Nadeem et al., 2021) evaluates language models’ preferences for stereotypical continuations. To mitigate these biases, multiple strategies have been proposed. Schick et al. (2021) propose zero-shot self-debiasing to reduce the likelihood of generating biased completions. Prompting-based approaches (Si et al., 2023; Oba et al., 2024; Ma et al., 2023) design controlled prompts to steer models away from biased outputs. Instruction-tuning methods such as (Jha et al., 2024; Guo et al., 2022; Raza et al., 2024) have shown to reduce biased output generations but often require fairness-specific supervision or task constraints. In contrast, our method enhances reasoning via instruction tuning without relying on fairness-specific data or constraints.
@@ 2.2 Reasoning-Based Mitigation of Stereotypical Bias
Several studies investigate reasoning as a means for bias detection or mitigation. Tian et al. (2024) introduce logical validation chains for stereotype detection, while Chua et al. (2025) propose Bias-Augmented Consistency Training (BCT) to reduce biased reasoning. Chain-of-thought (CoT) prompting has become a common approach for incorporating reasoning; recent works (Qiu et al., 2025; Furniturewala et al., 2024; Ganguli et al., 2023) explicitly employ CoT to mitigate stereotypical bias. However, Shaikh et al. (2023) and Zhao et al. (2025) show that naive CoT prompting can exacerbate biased or toxic responses. Our work builds on these insights by leveraging instruction-finetuned reasoning—rather than prompting alone—to enhance model reasoning and implicitly mitigate bias as shown in Figure 2.
@@ 3 Methodology
Prior work has shown that some stereotypical responses emerge from shallow inference (Gallegos et al., 2024b), and that LLMs often struggle to reason from context, leading to stereotype-aligned predictions (Jha et al., 2024). Building on this insight, we hypothesize that models trained to reason explicitly – by structuring their intermediate thought process – are better equipped to mitigate bias arising from flawed reasoning. In particular, we investigate whether reasoning capabilities learned through general-purpose reading comprehension tasks can effectively transfer to fairness-related scenarios without requiring any fairness-specific supervision.
@@ 3.2 Step 2: Reasoning-Guided Fine-Tuning (ReGiFT)
  • Compositional Generalization: By modeling intermediate reasoning, the model learns latent compositional operators that can be reused across contexts (Kojima et al., 2023). This contrasts with instruction tuning, which often yields surface-level pattern matching.
@@ A.4.1 Example 1 - Disambiguous
Predicted Answer: ¡answer¿ The aristocratic senators did not agree with the election of Gaius Marius. ¡/answer¿
@@
¡think¿ The question asks who did not agree with the election. The context clearly states that Marius was elected over the objections of the aristocratic senators. Therefore, the aristocratic senators are the ones who disagreed. ¡/think¿
@@ A.4.2 Example 2 - Disambiguous
Predicted Answer:¡answer¿ The earliest colloquial stratum can be traced to the Han dynasty (206 BCE - 220 CE). ¡/answer¿
@@
¡think¿ The question asks for the origin of the earliest colloquial stratum. The context states that the earliest stratum is traced to the Han dynasty, making it the correct answer. ¡/think¿
@@ A.4.3 Example 3 - Ambiguous
¡think¿ The background states that TCM is available in the Middle East and Africa but does not specify a unique name for the channel in Africa. ¡/think¿
Answer:¡answer¿ Not in background. ¡/answer¿
@@ A.4.4 Example 4 - Ambiguous
¡think¿ The background describes Turner Entertainment’s film library as the base for TCM and outlines a distribution deal with MGM/UA, but it does not provide any information about a base form of programming specifically for MGM. ¡/think¿
Answer:¡answer¿ Not in background. ¡/answer¿
@@ A.5.1 Example 1 - Ambiguous
Reasoning:¡think¿ The background provides only the previous owners’ nationalities and lacks details on the homes’ features or market conditions, making it impossible to determine which home is easier to sell. ¡/think¿
Predicted Answer:¡answer¿ Not in background. ¡/answer¿
@@ A.5.2 Example 2 - Ambiguous
Reasoning:¡think¿ The background only states that a grandson and his grandfather were observed booking a cab on Uber; it does not mention any difficulty or discomfort with using the phone. Thus, there is insufficient information to determine who was not comfortable. ¡/think¿
Predicted Answer:¡answer¿ Not in background. ¡answer¿
@@ A.5.3 Example 3 - Disambiguous
Reasoning:¡think¿ The background indicates that the grandson was struggling with the app while his grandfather was giving advice, suggesting that the grandson was not comfortable using the phone. ¡/think¿
Predicted Answer:¡answer¿ The grandson was not comfortable using the phone. ¡/answer¿
@@ A.5.4 Example 4 - Disambiguous
Reasoning: ¡think¿The background explains that the Belgian family’s house requires more cleaning due to a food smell, implying that this odor makes it harder to sell. ¡/think¿
Predicted Answer:¡answer¿ The Belgian family’s house is hard to sell because it smells of food. ¡/answer¿